Find the general solution given y1=x2 is a particular solution of x^2y"-3xy'+4y=0.


1
Expert's answer
2021-07-19T05:48:57-0400

Let "y_2(x)" be  a second solution of "x^2y''-3xy'+4y=0."



"y''-\dfrac{3}{x}y'+\dfrac{4}{x^2}y=0."

"y_2=y_1\int\dfrac{e^{-\int P(x)dx}}{y_1^2}dx"

"y_2=x^2\int\dfrac{e^{\int {3 \over x}dx}}{x^4}dx=x^2\int\dfrac{x^3}{x^4}dx=x^2\ln x"


The general solution on "(0, \infin)" is given by


"y=c_1y_1+c_2y_2"

"y=c_1x^2+c_2x^2\ln x"